# Towards Better Chain-of-Thought: A Reflection on Effectiveness and Faithfulness

### v2

@@ Abstract
Jiachun Li1,2, Pengfei Cao1,2, Yubo Chen1,2 Jiexin Xu3, Huaijun Li3, Xiaojian Jiang3, Kang Liu1,2, Jun Zhao1,2 1School of Artificial Intelligence, University of Chinese Academy of Sciences 2The Key Laboratory of Cognition and Decision Intelligence for Complex Systems, Institute of Automation, Chinese Academy of Sciences 3China Merchants Bank {jiachun.li, pengfei.cao, yubo.chen, kliu, jzhao}@nlpr.ia.ac.cn
@@ 1 Introduction
Recently, with chain-of-thought (CoT) techniques (Wei et al., 2022), large language models (LLMs) are able to reason on complex tasks (Wang et al., 2023; OpenAI, 2023). By scaling the CoT process using reinforcement learning (RL), LLMs can even surpass human performance in competition-level mathematical problems (OpenAI, 2024; DeepSeek-AI et al., 2025). However, despite the significant success of the CoT, some studies find that it demonstrates poor performance on certain tasks (Sprague et al., 2024; Xu and Ma, 2024; Turpin et al., 2023; Lanham et al., 2023). In some cases, using CoT for the model’s reasoning is unnecessary or even harmful (Wang et al., 2024b; Li et al., 2024).
@@
In summary, our key contributions are as follows: (1) We identify key factors that influence CoT’s effectiveness on different reasoning tasks, including problem difficulty, information gain, and information flow. (2) We interpret the unfaithful CoT issue by jointly analyzing the information interaction among question, CoT, and answer. Based on experimental results, we demonstrate that the reason is that LLMs retrieve correct information (lost in the CoT) directly from the question when predicting answers. (3) As an application of our findings, we design a new method called QUIRE, which effectively improves the CoT’s performance from the effectiveness (up to 2.4% improvement) and faithfulness (up to 5.6% improvement). This indicates that enhancing CoT faithfulness can lead to an improvement in CoT effectiveness. We release the source code in the attached software package.
@@ 2.1 Chain-of-Thought Effectiveness
Since the emergence of CoT, a series of CoT-like approaches have further improved the model’s reasoning accuracy through various prompt designs (Wang et al., 2023; Madaan et al., 2023b; Zhou et al., 2023). Recently, the emergence of reasoning models such as DeepSeek-R1 (DeepSeek-AI et al., 2025) and o1 (OpenAI, 2024) has once again proven that CoT is highly effective in solving complex reasoning tasks such as mathematics and coding (Qi et al., 2024; Snell et al., 2024; Zeng et al., 2024; Lightman et al., 2024). However, another series of works shows that the effectiveness of CoT has limitations (Wang et al., 2024b; Xu and Ma, 2024). They demonstrate that CoT brings only limited improvements in knowledge and commonsense reasoning tasks (Sprague et al., 2024), and may even harm the model’s original performance (Li et al., 2024). Building on these studies, our work further investigates the key factors that control CoT’s effectiveness across different tasks.
@@ 3.4 Information Flow
Following previous works (Wu et al., 2023; Wang et al., 2024a; Li et al., 2024), we employ integrated gradient attribution (IGA) (Sundararajan et al., 2017) as our measuring method to capture the information flow between CoT and answer. Specifically, we first compute importance $I_{n,m}$ of input token $x_{n}$ to output token $y_{m}$ as follows:
@@ 5.1 Our Method
As mentioned in Cl.6, when unfaithful issues occur, LLMs maintain a strong causal relevance with the correct statement in the context during the answer prediction. Thus, here we first generate a raw answer $A$ with the SC method, then recall extra information by selecting the top-k context statements with the highest $AAE(S,A)$ (as marked with red in Figure 11). After recalling extra information, we incorporate these statements as additional hints into the input prompt, enabling the model to pay more attention to this information during the CoT generation.
@@ Appendix E Implementation Details of the Main Experiment
Here we provide a detailed account of the implementation specifics from the main experiments in §5. For SC, we generate 3 samples for each question since our method is also set to 3 paths. For our method, we recall top-3 statements in AAE recall and generate one CoT for each enhanced prompt. We release all the prompts we use in the attached software package.

### v3

@@ Abstract
Jiachun Li1,2, Pengfei Cao1,2111Corresponding authors., Yubo Chen1,2 Jiexin Xu3, Huaijun Li3, Xiaojian Jiang3, Kang Liu1,2, Jun Zhao1,2111Corresponding authors. 1School of Artificial Intelligence, University of Chinese Academy of Sciences 2The Key Laboratory of Cognition and Decision Intelligence for Complex Systems, Institute of Automation, Chinese Academy of Sciences 3China Merchants Bank {jiachun.li, pengfei.cao, yubo.chen, kliu, jzhao}@nlpr.ia.ac.cn
@@ 1 Introduction
Recently, with chain-of-thought (CoT) techniques (Wei et al., 2022), large language models (LLMs) are able to reason on complex tasks (Wang et al., 2023; OpenAI, 2023; Jin et al., 2024b; Li et al., 2024b). By scaling the CoT process using reinforcement learning (RL), LLMs can even surpass human performance in competition-level mathematical problems (OpenAI, 2024; DeepSeek-AI et al., 2025). However, despite the significant success of the CoT, some studies find that it demonstrates poor performance on certain tasks (Sprague et al., 2024; Xu and Ma, 2024; Turpin et al., 2023; Lanham et al., 2023). In some cases, using CoT for the model’s reasoning is unnecessary or even harmful (Wang et al., 2024b; Li et al., 2024c).
@@
In summary, our key contributions are as follows: (1) We identify key factors that influence CoT’s effectiveness on different reasoning tasks, including problem difficulty, information gain, and information flow. (2) We interpret the unfaithful CoT issue by jointly analyzing the information interaction among question, CoT, and answer. Based on experimental results, we demonstrate that the reason is that LLMs retrieve correct information (lost in the CoT) directly from the question when predicting answers. (3) As an application of our findings, we design a new method called QUIRE, which effectively improves the CoT’s performance from the effectiveness (up to 2.4% improvement) and faithfulness (up to 5.6% improvement). This indicates that enhancing CoT faithfulness can lead to an improvement in CoT effectiveness. Our code is available at: https://github.com/BugMakerzzz/better_cot.
@@ 2.1 Chain-of-Thought Effectiveness
Since the emergence of CoT, a series of CoT-like approaches have further improved the model’s reasoning accuracy through various prompt designs (Wang et al., 2023; Madaan et al., 2023b; Zhou et al., 2023). Recently, the emergence of reasoning models such as DeepSeek-R1 (DeepSeek-AI et al., 2025) and o1 (OpenAI, 2024) has once again proven that CoT is highly effective in solving complex reasoning tasks such as mathematics and coding (Qi et al., 2024; Snell et al., 2024; Zeng et al., 2024; Lightman et al., 2024). However, another series of works shows that the effectiveness of CoT has limitations (Wang et al., 2024b; Xu and Ma, 2024; Li et al., 2024a). They demonstrate that CoT brings only limited improvements in knowledge and commonsense reasoning tasks (Sprague et al., 2024), and may even harm the model’s original performance (Li et al., 2024c). Building on these studies, our work further investigates the key factors that control CoT’s effectiveness across different tasks.
@@ 3.4 Information Flow
Following previous works (Wu et al., 2023; Wang et al., 2024a; Li et al., 2024c; Jin et al., 2024a), we employ integrated gradient attribution (IGA) (Sundararajan et al., 2017) as our measuring method to capture the information flow between CoT and answer. Specifically, we first compute importance $I_{n,m}$ of input token $x_{n}$ to output token $y_{m}$ as follows:
@@ 5.1 Our Method
As mentioned in Cl.6, when unfaithful issues occur, LLMs maintain a strong causal relevance with the correct statement in the context during the answer prediction. Thus, here we first generate a raw answer $A$ with the Self-Consistency (SC) method, then recall extra information by selecting the top-k context statements with the highest $AAE(S,A)$ (as marked with red in Figure 11). After recalling extra information, we incorporate these statements as additional hints into the input prompt, enabling the model to pay more attention to this information during the CoT generation.
@@ 6 Conclusion
Acknowledgments
This work is supported by the National Natural Science Foundation of China (No. U24A20335, No. 62176257, No. 62406321). This work is also supported by the Youth Innovation Promotion Association CAS and the China Postdoctoral Science Foundation under Grant Number 2024M753500.
@@ Appendix E Implementation Details of the Main Experiment
Here we provide a detailed account of the implementation specifics from the main experiments in §5. For SC, we generate 3 samples for each question since our method is also set to 3 paths. For our method, we recall top-3 statements in AAE recall and generate one CoT for each enhanced prompt.